# Towards Proof Repair in Cubical Agda

### 02 Feb 2024 v2

@@ 4.1. Adding, Fast and Slow
Notice that the call to $\mathbb{Z}$ in the first term is replaced with one to depElimGZ in the second term. We also see that, in the course of this repair, we had to repair two other functions: +pos and +negsuc. These terms are repaired following the same algorithm, and are omitted for space.
We can also repair theorems about our types. In $\mathbb{Z}$, we are able to prove the following theorem, which states that 00 is a left identity for addition:
  $\mathbb{Z}$ : (z : $\mathbb{Z}$) normal-$\to$ z $\equiv$ ($\mathbb{Z}$Pos 0) $\mathbb{Z}$ z
By manually following the repair algorithm from Section 3, we in turn get a proof of the repaired theorem for GZ:
  add0LGZ : (z : GZ) normal-$\to$ z $\equiv$ (depConstrGZPos 0) +$GZ$ z
@@
Here, this manifests as follows: The most general proof that we repair from $\mathbb{Z}$ to GZ is the eliminator for $\mathbb{Z}$, which we use to define both addition and the proof about it. This gives us a repaired eliminator for GZ, which occurs inside of the repaired addition function and proof over GZ.
But this repaired eliminator is slow, and so are our repaired functions and proofs that use it. It is not the eliminator that one would naturally define for GZ. Rather, it internally computes a canonical representative of the equivalence class of that element, either of the form $(n,0)$ for $n\geq 0$ or $(0,n)$ for $n\geq 1$. Because these canonical elements operate like pos n and negsuc n from the original type, we can compute on them in the same way as we did the original type—in other words, it follows the inductive structure of $\mathbb{Z}$. This means that, for any computation, we must compute canonical elements of equivalence classes, which is wasteful.
Instead, we would much rather use the more natural function $\lambda(n_{1},n_{2})(m_{1},m_{2}).(n_{1}+m_{1},n_{2}+m_{2})$. We are able to define this in Cubical Agda by way of the term in Figure 7.
@@
addResp : (a a b b : $\mathbb{N}\times\mathbb{N}$) normal-$\to$ $\sim$ a a normal-$\to$ $\sim$ b b normal-$\to$ $\sim$ (addHelpFunc a b) (addHelpFunc a b’)
-- proof of addResp omitted
@@
  setQuotBinOp isReflR isReflR addHelpFunc addResp
Figure 7. Our fast addition function on the repaired integers.
@@
  addEqual = funExt
    ($\lambda$ x normal-$\to$ funExt ($\lambda$ y normal-$\to$ addEqualOnInputs x y))
@@
  add’0LGZ: (z: GZ) normal-$\to$
    z $\equiv$ addGZ (depConstrGZPos 0) z
@@
This points to an advantage over working in Coq, whose type theory has neither functional extensionality nor transport, forcing prior work in proof repair to take a more ad hoc approach to moving between slow and fast repaired functions. We discuss this and other takeaways in detail in Section 5.
@@ 4.2. Variations on a Theme of Queues
Next, we repair functions and proofs across a change in implementation of a queue data structure. This is motivated by an example from Angiuli et al. (2021), which showed that quotient types can be use to adjust certain relations more general than equivalences into equivalences for use with transport in Cubical Agda. That class of changes was cited in the Pumpkin Pi paper as an example that could not be expressed naturally in Coq with the original framework. In Cubical Agda, we can express this use case naturally, highlighting the benefits of working in a type theory with quotient types—something we discuss in detail in Section 5.1.1.
@@
~ : (List A $\times$ List A) normal-$\to$ (List A $\times$ List A) normal-$\to$ Type
~ (l1, l2) (l3, l4) = l1 ++ (rev l2) = l3 ++ (rev l4)
@@
Figure 8. One list queues (OLQ) and two list queues (TLQ).
@@
To do repair between these types, we need them to be isomorphic. However, List A and (List A $\times$ List A) are not naturally isomorphic. Based on our description, we would say that a two list queue (l1 , l2) corresponds to the queue l1 ++ (rev l2), where ++ is the list append operator. However, this is not an injective map; multiple two list queues correspond to a single one list queue. To resolve this, we will take a quotient of our type of two list queues. We define the equivalence relation (l1 , l2) $\sim$ (l3 , l4) $\iff$ l1 ++ (rev l2) = l3 ++ (rev l4) and quotient (List A $\times$ List A) by it, obtaining the type TLQ of two list queues seen in Figure 8. The resulting two types have an isomorphism by the function [(l1 , l2)] $\mapsto$ (l1 ++ (rev l2)), which is well defined on our quotient type and has as an inverse the function l $\mapsto$ [(l , [])].
@@
We wish to conduct repair from our one list queue type to our two list queue type. We first write the configuration for our isomorphism (Figure 9). We have defined the standard functions enqueue and dequeue defined for one list queues, and we wish to repair them to our newly defined two list queue type. We also have theorems about these functions whose proofs we want to repair. The main theorem states that enqueue and dequeue are related in the way we expect:
  returnOrEnq : A normal-$\to$ Maybe (Q $\times$ A) normal-$\to$ Q $\times$ A
  returnOrEnq a = Cubical.Data.Maybe.rec

### 15 Mar 2024 v3

@@ 6.1. Adding, Fast and Slow
Definition addGZ (z1 z2 : GZ) : GZ :=
  depRecGZ
    GZ
    (fun p : nat => add_posGZ z1 p)
    (fun p : nat => add_negsucGZ z1 p)
    z2.
Figure 9. The repaired term for addition from $\mathbb{Z}$ onto $\mathbb{Z}$, in Cubical Agda (left) and Coq (right). The repaired Cubical Agda term is manually derived, while the repaired Coq term is automatically generated.
Then, we can follow the repair algorithm to produce the repaired terms shown in Figure 9. The repair of the Cubical Agda term was performed manually. Notice that, in the Cubical Agda code, the call to $\mathbb{Z}$ in the first term is replaced with one to depElimGZ in the second term. On the other hand, repair of the Coq term was done automatically by our automation prototype by running the following command:
    Lift Z GZ in addZ as addGZ.
The transformation still behaves similarly. In the Coq code, the call to depRecZ is directly replaced with one to depRecGZ. We also see that, in the course of this repair, we had to repair two other functions: +pos and +negsuc in Cubical Agda, and their analogues add_posZ and add_negsucZ in Coq. These terms are repaired following the same algorithm, and are omitted for space.
It is worth noting that, while the functions in Coq are automatically repaired, our automation prototype does not currently attempt to generate proofs that the repaired functions are proper. Thus, the user is currently responsible for writing these proofs, though a future version of the automation could generate many of these proofs using proof search, such as Coq’s built in solve_proper tactic. Of course, for the Cubical Agda terms, this is not an issue, since it is not possible to define a function on a quotient type which is not proper.
$\mathbb{Z}$ : (z : $\mathbb{Z}$) normal-$\to$
  z $\equiv$ ($\mathbb{Z}$Pos 0) $\mathbb{Z}$ z
add0LGZ : (z : GZ) normal-$\to$
  z $\equiv$ (depConstrGZPos 0) +GZ z
add0LZ :  forall z : Z,
  z = addZ (depConstrZPos 0) z
add0LGZ : forall z : GZ,
  eq_GZ z (addGZ (depConstrGZPos 0) z)
Figure 10. A theorem whose proof we repaired, in Cubical Agda (left) and Coq (right). The Cubical Agda proof was manually repaired, while the Coq proof was automatically repaired.
We can also repair proofs. In Figure 10, we see old and new versions of a theorem whose proof we repaired (the repaired proofs are in the supplementary material). The theorem states that 00 is a left identity for addition. Again, we can see elements of one side of the configuration swapped for the other. However, we also see that in Coq, equality has been replaced with an equivalence relation on the type, reflecting the fact that we are making a claim about setoid equality.
As before, the repair on Cubical Agda terms was manual, and was automated for Coq. However, the current version of Pumpkin Pi does not support configurations with multiple eliminators, and proving add0LZ requires using both depRecZ and depElimPropZ. We circumvent this by first defining constants referring to all subterms which contain depRecZ, and then separately repairing those constants. Then, we rewrite the proof of the theorem using these constants and depElimPropZ. We then reconfigure Pumpkin Pi to use depElimPropZ (by calling Pumpkin Pi’s configuration command a second time), and then repair the rewritten theorem after reconfiguring. This is not a fundamental limitation of our approach, and is inherited from Pumpkin Pi; future versions can remove the need for this workaround by supporting multiple eliminators in a single configuration.
In addition to the mentioned terms, we also repair the successor and predecessor functions on integers, functions which add $n$ or subtract $n+1$ from an integer (which were used above to define general addition), and the proof of a theorem saying that 00 is a right identity for addition. In Coq, all these functions and proofs could be repaired automatically, while repair was conducted manually in Cubical Agda. All of these functions and proofs, as well as their repaired versions, can be found in the supplementary material.
@@
Here, this manifests as follows: The most general proof that we repair from $\mathbb{Z}$ to GZ is the eliminator for $\mathbb{Z}$, which we use to define both addition and the proof about it. This gives us a repaired eliminator for GZ, which occurs inside of the repaired addition function and proof over GZ. But this repaired eliminator is slow, and so are our repaired functions and proofs that use it. It is not the eliminator that one would naturally define for GZ. Rather, it internally computes a canonical representative of the equivalence class of that element, either of the form $(n,0)$ for $n\geq 0$ or $(0,n)$ for $n\geq 1$. Because these canonical elements operate like pos n and negsuc n from the original type, we can compute on them in the same way as we did the original type—in other words, it follows the inductive structure of $\mathbb{Z}$. This means that, for any computation, we must compute canonical elements of equivalence classes, which is wasteful.
Instead, we would much rather use the more natural function $\lambda(n_{1},n_{2})(m_{1},m_{2}).(n_{1}+m_{1},n_{2}+m_{2})$. We are able to define this by way of the term in Figure 11.
@@
[…]
addResp : (a a b b : $\mathbb{N}\times\mathbb{N}$) normal-$\to$
  $\sim$ a a normal-$\to$
  $\sim$ b b normal-$\to$
  $\sim$ (addHelpFunc a b) (addHelpFunc a b’)
-- proof of addResp omitted for space
@@
  setQuotBinOp
    isReflR
    isReflR
    addHelpFunc
    addResp
Definition fastAddGZ (a b : GZ) :=
  match b with
  | (b1, b2) =>
    match a with
    | (a1, a2) => (a1 + b1, a2 + b2)
    end
  end.
Figure 11. Our fast addition function on the repaired integers, in Cubical Agda (left) and Coq (right). In addGZ’, setQuotBinOp is a specialized eliminator that allows us to easily define a binary function on set quotients. The definition of fast addition in Coq is shorter because we do not need to prove respectfulness when defining it. The use of pattern matching is acceptable because this function is neither to be repaired nor a product of repair.
@@
  addEqual = funExt ($\lambda$ x normal-$\to$ funExt ($\lambda$ y normal-$\to$ addEqualOnInputs x y))
@@
  add’0LGZ : (z: GZ) normal-$\to$ z $\equiv$ addGZ (depConstrGZPos 0) z
@@
We can move between slow and fast functions like this in Coq as well, but (consistently with prior work) we must take a more ad hoc approach, since Coq’s type theory has neither functional extensionality nor transport. Instead, we can define our fastAddGZ function, and prove the theorem:
  Theorem addEqualFastAdd : forall (a b : GZ), eq_GZ (addGZ a b) (fastAddGZ a b).
Then, as long as we can rewrite all applications of addGZ in a theorem statement, we can obtain the corresponding theorem about fastAddGZ. For example, we can translate add0LGZ into the theorem:
  Theorem fastAdd0LGZ : forall (z : GZ), eq_GZ z (fastAddGZ (depConstrGZPos 0) z).
using only one rewrite by addEqualFastAdd. If applications of addGZ were inside opaque terms, however, we may not be able to view all the applications of addGZ, and thus could not rewrite like this. Higher order functions are one potential source of this trouble; for instance, if the definition of List.map is opaque, and we pass addGZ as the function to map over a list, we would not be able to access the application sites of addGZ and thus could not rewrite by addEqualFastAdd. If we knew that fastAddGZ = addGZ, we could perform the rewrite anyway, which is one advantage Cubical Agda’s approach holds over Coq.
@@ 6.2. Variations on a Theme of Queues
Next, we repair functions and proofs across a change in implementation of a queue data structure. This is motivated by an example from Angiuli et al. (2021), which showed that quotient types can be use to adjust certain relations more general than equivalences into equivalences for use with transport in Cubical Agda. That class of changes was cited in the Pumpkin Pi paper as an example that could not be expressed naturally in Coq with the original framework. In Cubical Agda, we can express this use case naturally, highlighting the benefits of working in a type theory with quotient types. In Coq, with our extensions to Pumpkin Pi, we can express this using setoids. Our Cubical Agda and Coq proofs for this case study can be found in two_list_queue_equivalence.agda and two_list_queue_equivalence.v, respectively.
@@
~ : (List A $\times$ List A) normal-$\to$
      (List A $\times$ List A) normal-$\to$
     Type
~ (l1, l2) (l3, l4) =
  l1 ++ (rev l2) = l3 ++ (rev l4)
@@
Definition OLQ := list A.
Definition TLQ := list A * list A.
Definition insOrder (q : TLQ) :=
  match q with
  | (l1, l2) => l1 ++ rev l2
  end.
Definition eq_queue (q1 q2 : TLQ) :=
  insOrder q1 = insOrder q2.
Instance eq_queue_equiv : Equivalence eq_queue.
Instance TLQ_setoid : Setoid TLQ  :=
  {equiv := eq_queue ;
  setoid_equiv := eq_queue_equiv}.
Figure 12. One list queues (OLQ) and two list queues (TLQ) in Cubical Agda (left) and Coq (right). The Setoid instance is not explicitly needed, but is included for reader clarity.
@@
To do repair between these types, we need them to be isomorphic, but the types List A and (List A $\times$ List A) are not naturally isomorphic. Based on our description, we would say that a two list queue (l1 , l2) corresponds to the queue l1 ++ (rev l2), where ++ is the list append operator. However, this is not an injective map; multiple two list queues correspond to a single one list queue. In Cubical Agda, we resolve this by taking a quotient of our type of two list queues. We define the equivalence relation (l1 , l2) $\sim$ (l3 , l4) $\iff$ l1 ++ (rev l2) = l3 ++ (rev l4) and quotient (List A $\times$ List A) by it, obtaining the type TLQ of two list queues seen in Figure 12. The resulting two types have an isomorphism by the function [(l1 , l2)] $\mapsto$ (l1 ++ (rev l2)), which is well defined on our quotient type and has as an inverse the function l $\mapsto$ [(l , [])]. In Coq, we do not have quotient types, but instead work in the setoid on (list A * list A) with the previously mentioned equivalence relation.
@@
We wish to conduct repair from our one list queue type to our two list queue type. We first write the configuration for our isomorphism. The Cubical Agda configuration can be found in Figures 19 and  20, while the Coq configuration can be found in Figures 21 and  22, both of which are in the appendix. We have the standard functions enqueue and dequeue defined for one list queues, and we wish to repair them to our newly defined two list queue type. We also have theorems about these functions whose proofs we want to repair. The main theorem states that enqueue and dequeue are related in the way we expect (Figure 13).
  returnOrEnq : A normal-$\to$ Maybe (OLQ $\times$ A) normal-$\to$ OLQ $\times$ A
  returnOrEnq a =
    Cubical.Data.Maybe.rec